OpenStudy (anonymous):

int (x^(1/4))/(1+x^(1/2))

OpenStudy (chaise):

\[\int\limits \frac{x ^{\frac{1}{4}}}{1+x^{\frac{1}{2}}}\] Is this what you're asking? Do you just want an answer, or how to solve?

OpenStudy (anonymous):

yes that is what I am asking. I would like to know how to solve the problem.

OpenStudy (anonymous):

hmm, I feel that this looks a lot like, u / (1+u^2), where u = x^(1/4). Perhaps such a substitution would help?

OpenStudy (anonymous):

put \[u= x^{\frac 1 4} \] afer simplification, you get \[ 4 \int \frac { u^4}{u^2 +2}du \] and continue after that

OpenStudy (anonymous):

Got it now thank you.

OpenStudy (anonymous):

cool. and remember to mark the question as closed once you have solved it. thanks!

OpenStudy (anonymous):

done.

OpenStudy (anonymous):

Notice that \[ \frac{u^4}{u^2+1}=u^2-1 +\frac{1}{u^2+1} \] Now it is easy.

OpenStudy (anonymous):

After integrating and replacing you get \[ \frac{4 x^{3/4}}{3}-4 \sqrt[4]{x}+4 \tan ^{-1}\left(\sqrt[4]{x}\right) + C \]